40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Apr 1 23:05:58 2011, kiwamu, Update, Green Locking, noise budget noise_budget.png
    Reply  Wed Apr 6 02:41:01 2011, kiwamu, Update, Green Locking, noise budget : some more noise  noise_budget.png
Message ID: 4491     Entry time: Wed Apr 6 02:41:01 2011     In reply to: 4482
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: noise budget : some more noise  

It turned out that the dark noise from the beat PD and the shot noise on the beat PD was overestimated.

So I corrected them in the plot of the last noise budget (#4482).

Additionally I added the end laser error signal in the plot. Here is the latest plot.

noise_budget.png

 The end laser error spectrum is big enough to cover most of the frequency range.
 (although it was taken at a different time from the other curves.)

Quote from #4482

According to the plot what we should do are :

  * More accurate PD noise measurement

  * More accurate shot noise estimation

ELOG V3.1.3-